Can someone please help with the following:
A single die is rolled 20 times.
What is the probability that a six is rolled exactly once, if it is known that at least one six is rolled?


Answer by stanbon(75887)

Can someone please help with the following:
A single die is rolled 20 times.
What is the probability that a six is rolled exactly once, if it is known that at least one six is rolled? ------
P(one 6) = 1 - P(no 6 in 20) = 1 - (5/6)^20 = 1 - 0.0261 = 0.9739
-------------------------------
Cheers,
Stan H.
-------------------